OpenStudy (anonymous):

Ratios and Equations Jim spent $51on 3 CD's. At this rate, how much would 8 CD's cost?

OpenStudy (anonymous):

51 /3 = x / 8

OpenStudy (kuoministers):

\[\frac{ 8 }{ 3 } \times \]

OpenStudy (kuoministers):

51

OpenStudy (kuoministers):

and that gives you?

OpenStudy (anonymous):

not understanding

OpenStudy (anonymous):

which part?

OpenStudy (anonymous):

divide 51 by 3. that gives you the unit rate

OpenStudy (anonymous):

that is 17

OpenStudy (anonymous):

51/3 = 17

OpenStudy (anonymous):

If you then take 17 and multiply it by 8 you will get your answer.

OpenStudy (anonymous):

so its 51/3=17/8

OpenStudy (anonymous):

no

OpenStudy (anonymous):

it is 17 x 8

OpenStudy (anonymous):

136

OpenStudy (anonymous):

What does that give you?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Good for you!

OpenStudy (anonymous):

the book is giving me a weird way to do it

OpenStudy (anonymous):

I am sure. but if you break it down to a table and list the equivalent fractions or ratios, you will see I am right

OpenStudy (anonymous):

it is telling me to either multiply or divide with arrows from numerator to numerator and denominator to denominator

OpenStudy (anonymous):

Okay. Hold on a sec. I can help with that too

OpenStudy (anonymous):

my dad did it your way

OpenStudy (anonymous):

he calls it new math

OpenStudy (anonymous):

It would be 51 x 8 divided by 3

OpenStudy (kuoministers):

just follow jstdandy way... ITS the easiest possible way im pretty sure

OpenStudy (kuoministers):

the way in your book is just to make you understand the question easier

OpenStudy (anonymous):

This new math is fun isn't it? lol

OpenStudy (anonymous):

my dad is pulling out his hair

OpenStudy (anonymous):

I am sure. If you have more of these type of problems Here are the steps to follow

OpenStudy (anonymous):

thank you for your help

OpenStudy (anonymous):

i have a bunch

OpenStudy (anonymous):

Take your complete fraction and set it side by side with your incomplete fradtion

OpenStudy (anonymous):

51/3 = x/8

OpenStudy (anonymous):

If your incomplete fraction has only a denominator, multiply the incomplete fractions denominator by the complete fractions numerator and divide by the complete fractions denominator. That will be the missing numerator.

OpenStudy (anonymous):

If your incomplete fraction has only a numerator, multiply the incomplete numerator by the denominator of the complete fraction and divide by the denominator of the complete fraction. That will give you your missing denominator.

OpenStudy (anonymous):

im still lost in setting it up. what goes over what?

OpenStudy (anonymous):

sorry :(

OpenStudy (anonymous):

Story problems?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok. Take the known information and make it a fraction. Like your first one said $51 for 3 cd's = 51/3

OpenStudy (anonymous):

Jim spent $51 on 3 CD's. At this rate, how much would 8 CD's cost? 51/3 = x/8

OpenStudy (anonymous):

because you needed to find the cost for 8 cd's your second fraction is ?/8

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

You need to find the the number to replace the ?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Set the others up the same way. What you know in a fraction

OpenStudy (anonymous):

but the ? becomes....136?

OpenStudy (anonymous):

Yes. it is the cost so, this problem was cost over quantity cost/quantity Your other problems with have similar information

OpenStudy (anonymous):

so it would look like: 51/3=136/8

OpenStudy (anonymous):

x=17

OpenStudy (anonymous):

YES! nice job! I feel bad, but I will have to check back with you. I need run an errand.

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

no problem. thank you for your help

OpenStudy (anonymous):

Wait, no. x=136

OpenStudy (anonymous):

oh, ok :)))))))